How many years will it take for an initial investment of 10,000 dollars to grow to 29,000 dollars? Assume a rate of interest of 6% compounded continuously.
Answer by stanbon(75887)

How many years will it take for an initial investment of 10,000 dollars to grow to 29,000 dollars? Assume a rate of interest of 6% compounded continuously.
--------
A(t) = P*e^(rt)
----
29,000 = 10,000*e^(0.06t)
-------------
e^(0.06t) = 29
---
0.06t = ln(29)
----
t = 56 years
==================
Cheers,
Stan H.
==================